like a good economist, you calculate the opportunity cost of getting your college degree. suppose that at your university you will pay 12,000 each year for tuition, 2,000 each year for books, and 8,000 per year for room and board. before you left for college, your boss at your high school job offered you a job paying 20,000 per year.

Assume if you decided not college, your parents would not let live at home.

What is your opportunity cost for four years of college?


1
Expert's answer
2021-09-02T16:44:17-0400

Opportunity cost is the cost of the next best alternative. If the person is not attending college then he would save his tuition and textbook fees but would still have to pay for room and board expenses. He will also not able to earn $20000 earn per year from his job.

Therefore, opportunity cost =

"(12000 + 2000 + 20000) \times 4 = 34000 \times 4 = \$136000"